OpenStudy (anonymous):

lim(x----->0) {1/x}^x

OpenStudy (anonymous):

I don't know of this helps, but: \[\lim_{x\rightarrow0}\left(\dfrac{1}{x}\right)^x=\lim_{x\rightarrow0}\dfrac{1^x}{x^x}=\dfrac{\lim_{x\rightarrow0}1^x}{\lim_{x\rightarrow0}x^x}=\dfrac{1}{\lim_{x\rightarrow0}x^x}\]

OpenStudy (anonymous):

but 0^0 doesn't make any sense

OpenStudy (anonymous):

\[\large\begin{align*} \lim_{x\to0}\left(\frac{1}{x}\right)^x&=\lim_{x\to0^+}e^{\ln\left(\frac{1}{x}\right)^x}\\\\ &=\lim_{x\to0}e^{x\ln\frac{1}{x}}\\\\ &=e^{{\displaystyle\lim_{x\to0^+}}~x\ln\frac{1}{x}}\\\\ &=e^{{\displaystyle\lim_{x\to0^+}}~\dfrac{\ln\frac{1}{x}}{\frac{1}{x}}}\\\\ &=e^{{\displaystyle\lim_{x\to0^+}}~\dfrac{-\ln x}{\frac{1}{x}}}\\\\ &=e^{-{\displaystyle\lim_{x\to0^+}}~\dfrac{\ln x}{\frac{1}{x}}}\\\\ &=e^{-{\displaystyle\lim_{x\to0^+}}~\dfrac{\frac{1}{x}}{-\frac{1}{x^2}}}&\text{via L'Hopital's rule}\\\\ &=e^{{\displaystyle\lim_{x\to0^+}}~\dfrac{1}{\frac{1}{x}}}\\\\ &=e^{{\displaystyle\lim_{x\to0^+}}x}\\\\ &=e^0\\\\ &=1 \end{align*}\]

OpenStudy (anonymous):

yeah it looks right

geerky42 (geerky42):

@micahwood50 kind of does help. I'm finish where he stopped. Let's focus on limit only, assume it's one-sided. \[\begin{align*} \lim_{x\rightarrow0^+} x^x&= \lim_{x\rightarrow0^+} e^{\ln(x^x)} \\~\\ &=\lim_{x\rightarrow0^+}e^{x\ln(x)}\\~\\ &=e^{~\displaystyle\lim_{x\rightarrow0^+}x\ln (x)}~~~\text{This will get us form 0/}\infty,\\&~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\text{ so let move x to denominator as 1/x}\\~\\ &=e^{~\displaystyle\lim_{x\rightarrow0^+}\frac{\ln (x)}{\frac{1}{x}}}~~~ \text{The form is now 0/0, so let's use L'Hospital's rule}\\~\\ &=e^{~\displaystyle\lim_{x\rightarrow0^+}\frac{\frac{1}{x}}{-\frac{1}{x^2}}}\\~\\ &=e^{~\displaystyle\lim_{x\rightarrow0^+}\frac{-x^2}{x}}\\~\\ &=e^{~\displaystyle\lim_{x\rightarrow0^+}-x}\\~\\ &=e^0\\~\\ &=\boxed{1} \end{align*}\] So we have \(\dfrac{1}{\displaystyle \lim_{x\rightarrow0^+}x^x}=\dfrac{1}{1}=\boxed{1}\)

geerky42 (geerky42):

Though I am confused about how you can switch from two-sided limit to one sided limit... Can you explain? @SithsAndGiggles

OpenStudy (anonymous):

@geerky42, the switch to a one-sided limit has to do with the introduction of the logarithm. \(\ln\dfrac{1}{x}\) is only defined for \(\dfrac{1}{x}>0\), i.e. \(x>0\) \(\bigg(\)since all numbers \(\dfrac{1}{x}\) are positive, but \(\dfrac{1}{x}\) is undefined at \(x=0\bigg)\).